# On the Convergence of Hermitian Dynamic Mode Decomposition

### v1

@@ Abstract
In this work, we study the convergence of Hermitian Dynamic Mode Decomposition (DMD) to the spectral properties of self-adjoint Koopman operators. Hermitian DMD is a data-driven method for approximating the Koopman operator associated with an unknown nonlinear dynamical system from discrete-time snapshots, while preserving the self-adjointness of the operator on its finite-dimensional approximations. We show that, under suitable conditions, the eigenvalues and eigenfunctions of HDMD converge to the spectral properties of the underlying Koopman operator. Along the way, we establish a general theorem on the convergence of spectral measures, and demonstrate our results numerically on the two-dimensional Schrödinger equation.
@@ 1 Introduction
where $\mathbf{x}\in\Omega$ denotes the state of the system, and $\Omega\subseteq\mathbb{R}^{d}$ is the state-space for $d\geq 1$. The (typically) nonlinear function $\mathbf{F}:\Omega\rightarrow\Omega$ governs the system’s evolution and is unknown. We assume that our knowledge of the system is limited to $M\geq 1$ discrete-time snapshots of the system, i.e., one has only access to a finite dataset of the form
@@
Data in the form of (1.2) can collected from either one long trajectory or multiple shorter trajectories, and acquired via experimental observations or numerical simulations. In general, one aims to exploit the snapshot data to infer and reconstruct properties of the underlying dynamical system given by (1.1). With the arrival of big data and machine learning, this data-driven viewpoint is currently undergoing a renaissance. Examples of applications of this framework arise naturally across many scientific fields, including fluid dynamics schmid2010dynamic , epidemiology rowley2009spectral , chemistry narasingam2019koopman , and neuroscience brunton2016extracting , to name a few.
One of the most prominent algorithms for data-driven analysis of dynamical systems is Dynamic Mode Decomposition (DMD), which is closely connected with Koopman operators. In 1931, Koopman introduced an operator-theoretic approach to dynamical systems, initially to describe Hamiltonian systems koopman1931hamiltonian . This theory was further expanded by Koopman and von Neumann koopman1932dynamical to include systems with continuous spectra. A Koopman operator $\mathcal{K}$ lifts a nonlinear system (1.1) into an infinite-dimensional space of observable functions $g:\Omega\rightarrow\mathbb{C}$ as
@@
Through this approach, the evolution dynamics become linear, enabling the use of generic solution techniques based on spectral decompositions. DMD was initially developed in the context of fluid dynamics schmid2010dynamic . Earlier, Mezić introduced the Koopman mode decomposition mezic2005spectral , providing a theoretical basis for Rowley et al. to connect DMD with Koopman operators rowley2009spectral . However, the standard DMD algorithm is based on linear observables and generally fails to capture truly nonlinear phenomena. To address this limitation, Extended DMD (EDMD) williams2015data extends the DMD algorithm to nonlinear observables. Then, under suitable conditions, in the large-data limit $M\rightarrow\infty$, EDMD converges to the numerical approximation obtained by a Galerkin method in the limit of large data sets. At its core, EDMD is a projection method that aims to compute spectral properties of Koopman operators.
@@
In a recent paper baddoo2023physics , Baddoo et al. provided a unified framework, denoted as Physics-Informed Dynamic Mode Decomposition (piDMD), for imposing physics constraints in the DMD framework with a linear choice of dictionary. For some constraints, this framework can be extended to the setting of EDMD, thus providing an approximation of Koopman operators in the spirit of geometric integration hairer2010geometric . A scenario where it is possible to extend piDMD to nonlinear observables occurs when the Koopman operator associated with the dynamical system is Hermitian, and one aims to compute finite-dimensional Hermitian approximations to preserve spectral properties. The corresponding algorithm, known as Hermitian Dynamic Mode Decomposition, has been studied extensively by Zlatko drmac2022lapack2 . However, there is currently no proof of convergence of this method to the spectral properties of the underlying Koopman operator, despite encouraging numerical results baddoo2023physics ; drmac2022lapack2 .
@@
This note addresses this gap by showing the convergence of Hermitian Dynamic Mode Decomposition to the spectral properties of self-adjoint Koopman operators. Along the way, we derive Theorem 4.5, a result that may be of independent interest in the wider spectral community. For methods that compute spectral measures of general self-adjoint systems, see colbrook2019computing ; colbrook2020computing . All of our results naturally extend to the case when the Koopman operator is skew-Hermitian, another structure of broad interest. As an example, Koopman generators of invertible measure-preserving continuous-time dynamical systems are skew-adjoint. Our results can also be extended to stochastic dynamical systems and the stochastic Koopman operator mezic2000comparison ; vcrnjaric2020koopman ; wanner2022robust ; colbrook2023beyond .
@@
It is worth pointing out the existence of a myriad of papers on Koopman operators and DMD, as evidenced by various surveys budivsic2012applied ; brunton2021modern ; colbrook2023multiverse ; mezic2013analysis ; otto2021koopman ; schmid2022dynamic . For example, the survey colbrook2023multiverse characterizes structure-preserving methods as one of the flavors of DMD. Despite this widespread interest, convergence results pertaining to the relevant spectral properties of Koopman operators remain decidedly rare. [^1] In particular, spectral approximations produced by EDMD do not generally converge. Exceptions to this rule include methods with theoretical guarantees, such as Hankel-DMD arbabi2017ergodic , Residual DMD colbrook2023residualJFM ; colbrook2021rigorous , Measure-Preserving EDMD colbrook2023mpedmd , compactification methods das2021reproducing ; valva2023consistent , and periodic approximation methods govindarajan2019approximation ; govindarajan2021approximation . While related to the present note, the latter three approaches assume that the system is measure-preserving, which differs from the setup addressed here. Given the keen interest in convergence results, we hope this note will encourage further exploration into the conditions under which methods like piDMD converge.
@@ 2.1 Koopman operators
The above Hilbert space setting is standard in most of the Koopman literature. It is crucial to recognize that the Koopman operator is not uniquely defined by the dynamical system in (1.1); rather, it is fundamentally dependent on the choice of the space of observables $\mathcal{F}$. Since $\mathcal{K}$ acts on an infinite-dimensional function space, we have exchanged the nonlinearity in (1.1) for an infinite-dimensional linear system. This means that the spectral properties of $\mathcal{K}$ can be significantly more complex than those of a finite matrix, making them more challenging to compute colbrook2020PhD ; ben2015can .
@@ 2.2 Extended dynamic mode decomposition
$\mathbf{\Psi}(\mathbf{x})=\begin{bmatrix}\psi_{1}(\mathbf{x})&\cdots&\psi_{{N}}(\mathbf{x})\end{bmatrix}\in\mathbb{C}^{1\times{N}},\quad\mathbf{x}\in\Omega.$
@@
where $\dagger$ denotes the Moore–Penrose pseudoinverse. The second equality follows since $\mathbf{W}^{1/2}\mathbf{\Psi}_{X}$ has linearly independent columns. Note that this solution is independent of the matrix $\mathbf{C}$. However, a suitable choice of $\mathbf{C}$ is vital once constraints on the matrix $\mathbf{K}$ are added to the optimization problem in (2.2colbrook2023mpedmd . In the case where the quadrature weights are equal and $\mathbf{\Psi}$ is the state (i.e., a linear dictionary), then $\mathbf{K}^{\top}$ is the transpose of the classical DMD matrix.
@@
Therefore, in the large data limit, $\mathbf{K}=\mathbf{G}^{\dagger}\mathbf{A}$ approaches a matrix representation of the operator $\mathcal{P}_{V_{N}}\mathcal{K}\mathcal{P}_{V_{N}}^{*}$, where $\mathcal{P}_{V_{N}}:L^{2}(\Omega,\omega)\rightarrow V_{N}$ denotes the orthogonal projection onto the subspace $V_{N}$. In essence, EDMD is a Galerkin method. The EDMD eigenvalues thus approach the spectrum of $\mathcal{P}_{V_{N}}\mathcal{K}\mathcal{P}_{V_{N}}^{*}$, and EDMD is an example of the so-called finite section method bottcher1983finite .
@@ 2.3 Spectral measures of self-adjoint Koopman operators
From now on, we assume that $\mathcal{K}$ is a self-adjoint operator acting on $L^{2}(\Omega,\omega)$. Under this condition, the spectral theorem [^3] For readers unfamiliar with the spectral theorem, halmos1963does provides an excellent and readable introduction. (conway2019course, , Thm. X.4.11) allows us to diagonalize the Koopman operator $\mathcal{K}$, and its spectrum $\mathrm{Sp}(\mathcal{K})$ lies within the real line $\mathbb{R}$ since $\mathcal{K}$ is self-adjoint. There is a projection-valued measure $\mathcal{E}$ supported on $\mathrm{Sp}(\mathcal{K})$, which associates an orthogonal projector with each Borel measurable subset of $\mathbb{R}$. For such a subset $S\subset\mathbb{R}$, $\mathcal{E}(S)$ is a projection onto the spectral elements of $\mathcal{K}$ inside $S$. For any observable $g\in\mathcal{D}(\mathcal{K})$,
@@
The spectral measure of $\mathcal{K}$ with respect to $g\in L^{2}(\Omega,\omega)$ is a signature for the forward-time dynamics of (1.1). In this paper, we show that the spectral measures of Hermitian DMD converge to the ones of $\mathcal{K}$.
@@ 3 Hermitian Dynamic Mode Decomposition
When the Koopman operator $\mathcal{K}$ is Hermitian, i.e., $\mathcal{K}=\mathcal{K}^{*}$, a natural constraint is to preserve the Hermitian property on its finite-dimensional approximation, $\mathbf{K}$. This ensures that the spectral properties of $\mathbf{K}$ are consistent with those of $\mathcal{K}$ as the size of the dictionary increases. Generally, the solution to (2.3) is not Hermitian, and diverse strategies have been proposed to enforce this constraint drmac2022lapack2 . Here, we consider the Hermitian DMD algorithm introduced in baddoo2023physics .
@@
$\langle\Psi\mathbf{g},\Psi\mathbf{h}\rangle=\sum_{j,k=1}^{N}\overline{h_{j}}g_{k}\langle\psi_{k},\psi_{j}\rangle\approx\sum_{j,k=1}^{N}\overline{h_{j}}g_{k}\mathbf{G}_{j,k}=\mathbf{h}^{*}\mathbf{G}\mathbf{g}.$ (3.1)
@@
Here, we recognize a symmetric Procrustes problem baddoo2023physics ; higham1988symmetric of the form
@@
for $1\leq i,j,\leq N$. Here, the coefficients $c_{ij}$ are the entries of the matrix $\mathbf{C}=\mathbf{U}^{*}\mathbf{Y}\mathbf{V}$. As observed by Higham higham1988symmetric , this algorithm is backward stable following the backward stability of the Golub–Reinsch SVD algorithm (golub1996matrix, , Sec. 5.5.8).
We make an additional observation compared to the original formulation in baddoo2023physics , which considerably simplifies the algorithm to solve (3.2). In the case of (3.2), we have $\mathbf{X}=\mathbf{W}^{1/2}\mathbf{\Psi}_{X}\mathbf{G}^{-1/2}$ and $\mathbf{Y}=\mathbf{W}^{1/2}\mathbf{\Psi}_{Y}\mathbf{G}^{-1/2}$, and the symmetric Procrustes algorithm can be simplified since $\mathbf{X}^{*}\mathbf{X}=\mathbf{I}$, the identity matrix. Hence, one can select $\boldsymbol{\Sigma}$ and $\mathbf{V}$ to be the identity matrices, and the solution becomes
@@
Therefore, if the convergence in (2.5) holds, the large data limit of Hermitian DMD becomes a Galerkin approximation of the Koopman operator since $\langle\mathcal{K}\psi_{k},\psi_{j}\rangle=\langle\psi_{k},\mathcal{K}\psi_{j}\rangle$. Additionally, the formula given by (3.3) is more computationally efficient than the standard symmetric Procrustes approach employed in baddoo2023physics as it avoids the computation of the SVD of $\mathbf{X}$, as well as a matrix square-root, which can be numerically unstable when $\mathbf{G}$ is ill-conditioned.
Given this convergence result, it is natural to study the convergence of this approximation as the size of the dictionary increases. It is well-known that spectra of Galerkin approximations of operators can suffer from discretization issues such as spectral pollution (spurious eigenvalues), spectral invisibility (missing parts of the spectrum), instabilities, and so forth colbrook2019compute ; colbrook4 ; colbrook3 . These issues occur and, in fact, can worsen as the dictionary’s size increases. Hence the convergence of the Hermitian DMD algorithm might be difficult to establish. Nevertheless, we will show in Section 4 that the (scalar-valued) spectral measures converge weakly, thus providing theoretical guarantees for Hermitian DMD. One crucial property exploited in the proof is the self-adjointness of the finite-dimensional approximations, as imposed by Hermitian DMD.

### v2

@@ Abstract
We study the convergence of Hermitian Dynamic Mode Decomposition (DMD) to the spectral properties of self-adjoint Koopman operators. Hermitian DMD is a data-driven method that approximates the Koopman operator associated with an unknown nonlinear dynamical system, using discrete-time snapshots. This approach preserves the self-adjointness of the operator in its finite-dimensional approximations. We prove that, under suitably broad conditions, the spectral measures corresponding to the eigenvalues and eigenfunctions computed by Hermitian DMD converge to those of the underlying Koopman operator. This result also applies to skew-Hermitian systems (after multiplication by $i$), applicable to generators of continuous-time measure-preserving systems. Along the way, we establish a general theorem on the convergence of spectral measures for finite sections of self-adjoint operators, including those that are unbounded, which is of independent interest to the wider spectral community. We numerically demonstrate our results by applying them to two-dimensional Schrödinger equations.
@@
journal: ArXiv
[…]
[Imperial] organization=Department of Mathematics, Imperial College London, city=London, postcode=SW7 2AZ, country=UK
@@ 1 Introduction
where $\mathbf{x}\in\Omega$ denotes the state of the system, and $\Omega\subseteq\mathbb{R}^{d}$ is the state-space for $d\in\mathbb{N}$. The (typically) nonlinear function $\mathbf{F}:\Omega\rightarrow\Omega$ governs the system’s evolution and is unknown. We assume that our knowledge of the system is limited to $M\geq 1$ discrete-time snapshots of the system, i.e., one has only access to a finite dataset of the form
@@
This snapshot data can be collected from either one long trajectory or multiple shorter trajectories and acquired via experimental observations or numerical simulations. In general, one aims to use this data to infer and reconstruct properties of the underlying dynamical system given by (1.1). With the arrival of big data and machine learning, this data-driven viewpoint is currently undergoing a renaissance. Examples of applications of this framework arise naturally across many scientific fields, including fluid dynamics [1], epidemiology [2], chemistry [3], and neuroscience [4], to name a few.
One of the most prominent algorithms for data-driven analysis of dynamical systems is Dynamic Mode Decomposition (DMD), which is closely connected with Koopman operators. In 1931, Koopman introduced an operator-theoretic approach to dynamical systems, initially to describe Hamiltonian systems [5]. This theory was further expanded by Koopman and von Neumann [6] to include systems with continuous spectra. A Koopman operator $\mathcal{K}$ lifts a nonlinear system (1.1) into an infinite-dimensional space of observable functions $g:\Omega\rightarrow\mathbb{C}$ as
@@
Through this approach, the evolution dynamics become linear, enabling the use of generic solution techniques based on spectral decompositions. DMD was initially developed in the context of fluid dynamics [7, 1]. Earlier, Mezić introduced the Koopman mode decomposition [8], providing a theoretical basis for Rowley et al. to connect DMD with Koopman operators [2]. However, the standard DMD algorithm is based on linear observables and generally fails to capture truly nonlinear phenomena. To address this limitation, Extended DMD (EDMD) [9] extends the DMD algorithm to nonlinear observables. Then, under suitable conditions, in the large-data limit $M\rightarrow\infty$, EDMD converges to the numerical approximation obtained by a Galerkin method in the limit of large data sets. At its core, EDMD is a projection method that aims to compute the spectral properties of Koopman operators. It is important to realize that EDMD only converges in the strong operator topology to $\mathcal{K}$ [10] [^1] Essentially, this means that the action of EDMD on observables converges to the action of the Koopman operator on observables. This form of convergence implies that limit points of eigenpairs of EDMD are also eigenpairs of the Koopman operator if the limiting vector is non-zero [10, Theorem 4]. As the authors of [10] highlight, this is an extremely weak form of convergence. and its spectral properties need not converge [11, Example 2]. For ways of overcoming this and classifications of how difficult Koopman spectral computations are, see [12].
@@
In a recent paper [13], Baddoo et al. provided a unified framework, Physics-Informed Dynamic Mode Decomposition (piDMD), for imposing physical constraints in DMD with a linear choice of dictionary. For some constraints, this framework can be extended to the setting of EDMD, thus providing an approximation of Koopman operators in the spirit of geometric integration [14]. A scenario where it is possible to extend piDMD to nonlinear observables occurs when the Koopman operator associated with the dynamical system is Hermitian, and one aims to compute finite-dimensional Hermitian approximations to preserve spectral properties. The corresponding algorithm, known as Hermitian Dynamic Mode Decomposition, has been studied extensively by Drmač [15]. However, there is currently no proof of convergence of this method to the spectral properties of the underlying Koopman operator, despite encouraging numerical results [13, 15].
@@
This note addresses this gap by showing the convergence of Hermitian Dynamic Mode Decomposition to the spectral properties of self-adjoint Koopman operators. Along the way, we derive Theorem 4.5, a result that may be of independent interest in the wider spectral community. For methods that compute spectral measures of general self-adjoint systems, see [16, 17]. All of our results naturally extend to the case when the Koopman operator is skew-Hermitian, another structure of broad interest. As an example, Koopman generators of invertible measure-preserving continuous-time dynamical systems are skew-adjoint. Hence, our analysis carries over to the application of (skew) Hermitian DMD to DMD methods that approximate such generators [18]. Our results can also be extended to stochastic dynamical systems and the stochastic Koopman operator [19, 20, 21, 22].
@@
It is worth pointing out the myriad of papers on Koopman operators and DMD, as evidenced by various surveys [23, 24, 25, 26, 27, 28]. For example, the survey [25] characterizes structure-preserving methods as one of the flavors of DMD. Despite this widespread interest, convergence results pertaining to the relevant spectral properties of Koopman operators remain decidedly rare. Exceptions to this rule include methods with theoretical guarantees, such as Hankel-DMD [29], Residual DMD [30, 31], Rigged DMD [32], Measure-Preserving EDMD [33], compactification methods [34, 35], and periodic approximation methods [36, 37]. While related to the present note, the latter four approaches assume that the system is measure-preserving, which differs from the setup addressed here. Given the keen interest in convergence results, we hope this note will encourage further exploration into the conditions under which methods like piDMD converge.
@@ 2.1 Koopman operators
The above Hilbert space setting is standard in most of the Koopman literature. It is crucial to recognize that the Koopman operator is not uniquely defined by the dynamical system in (1.1); rather, it is fundamentally dependent on the choice of the space of observables $\mathcal{F}$. Since $\mathcal{K}$ acts on an infinite-dimensional function space, we have exchanged the nonlinearity in (1.1) for an infinite-dimensional linear system. This means that the spectral properties of $\mathcal{K}$ can be significantly more complex than those of a finite matrix, making them more challenging to compute [38, 39].
@@ 2.2 Extended dynamic mode decomposition
$\mathbf{\Psi}(\mathbf{x})=\begin{bmatrix}\psi_{1}(\mathbf{x})&\cdots&\psi_{{N}}(\mathbf{x})\end{bmatrix}\in\mathbb{C}^{1\times{N}},\quad\mathbf{x}\in\Omega.$ (2.2)
@@
where $\dagger$ denotes the Moore–Penrose pseudoinverse. The second equality follows since $\mathbf{W}^{1/2}\mathbf{\Psi}_{X}$ has linearly independent columns. Note that this solution is independent of the matrix $\mathbf{C}$. However, a suitable choice of $\mathbf{C}$ is vital once constraints on the matrix $\mathbf{K}$ are added to the optimization problem in (2.3[33]. In the case where the quadrature weights are equal and $\mathbf{\Psi}$ is the state (i.e., a linear dictionary), then $\mathbf{K}^{\top}$ is the transpose of the classical DMD matrix.
@@
Therefore, in the large data limit, $\mathbf{K}=\mathbf{G}^{\dagger}\mathbf{A}$ approaches a matrix representation of the operator $\mathcal{P}_{V_{N}}\mathcal{K}\mathcal{P}_{V_{N}}^{*}$, where $\mathcal{P}_{V_{N}}:L^{2}(\Omega,\omega)\rightarrow V_{N}$ denotes the orthogonal projection onto the subspace $V_{N}$. In essence, EDMD is a Galerkin method. The EDMD eigenvalues thus approach the spectrum of $\mathcal{P}_{V_{N}}\mathcal{K}\mathcal{P}_{V_{N}}^{*}$, and EDMD is an example of the so-called finite section method [40].
@@ 2.3 Spectral measures of self-adjoint Koopman operators
From now on, we assume that $\mathcal{K}$ is a self-adjoint operator acting on $L^{2}(\Omega,\omega)$. Under this condition, the spectral theorem [^3] For readers unfamiliar with the spectral theorem, [41] provides an excellent and readable introduction. [42, Thm. X.4.11] allows us to diagonalize the Koopman operator $\mathcal{K}$, and its spectrum $\mathrm{Sp}(\mathcal{K})$ lies within the real line $\mathbb{R}$ since $\mathcal{K}$ is self-adjoint. There is a projection-valued measure $\mathcal{E}$ supported on $\mathrm{Sp}(\mathcal{K})$, which associates an orthogonal projector with each Borel measurable subset of $\mathbb{R}$. For such a subset $S\subset\mathbb{R}$, $\mathcal{E}(S)$ is a projection onto the spectral elements of $\mathcal{K}$ inside $S$. For any observable $g\in\mathcal{D}(\mathcal{K})$,
@@
The spectral measure of $\mathcal{K}$ with respect to $g\in L^{2}(\Omega,\omega)$ is a signature for the forward-time dynamics of (1.1). In this paper, we show that the spectral measures computed by Hermitian DMD converge to those of $\mathcal{K}$.
@@ 3 Hermitian Dynamic Mode Decomposition
When the Koopman operator $\mathcal{K}$ is Hermitian, i.e., $\mathcal{K}=\mathcal{K}^{*}$, a natural constraint is to preserve the Hermitian property on its finite-dimensional approximation, $\mathbf{K}$. This ensures that the spectral properties of $\mathbf{K}$ are consistent with those of $\mathcal{K}$ as the size of the dictionary increases. Generally, the solution to (2.4) is not Hermitian, and diverse strategies have been proposed to enforce this constraint [15]. Here, we consider the Hermitian DMD algorithm introduced in [13].
@@
$\langle\mathbf{\Psi}\mathbf{g},\mathbf{\Psi}\mathbf{h}\rangle=\sum_{j,k=1}^{N}\overline{h_{j}}g_{k}\langle\psi_{k},\psi_{j}\rangle\approx\sum_{j,k=1}^{N}\overline{h_{j}}g_{k}\mathbf{G}_{j,k}=\mathbf{h}^{*}\mathbf{G}\mathbf{g}.$ (3.1)
@@
Here, we recognize a symmetric Procrustes problem [13, 43] of the form
@@
for $1\leq i,j,\leq N$. Here, the coefficients $c_{ij}$ are the entries of the matrix $\mathbf{C}=\mathbf{U}^{*}\mathbf{Y}\mathbf{V}$. As observed by Higham [43], this algorithm is backward stable following the backward stability of the Golub–Reinsch SVD algorithm [44, Sec. 5.5.8].
We make an additional observation compared to the original formulation in [13], which considerably simplifies the algorithm to solve (3.2). In the case of (3.2), we have $\mathbf{X}=\mathbf{W}^{1/2}\mathbf{\Psi}_{X}\mathbf{G}^{-1/2}$ and $\mathbf{Y}=\mathbf{W}^{1/2}\mathbf{\Psi}_{Y}\mathbf{G}^{-1/2}$, and the symmetric Procrustes algorithm can be simplified since $\mathbf{X}^{*}\mathbf{X}=\mathbf{I}$, the identity matrix. Hence, one can select $\boldsymbol{\Sigma}$ and $\mathbf{V}$ to be the identity matrices, and the solution becomes
@@
Therefore, if the convergence in (2.6) holds, the large data limit of Hermitian DMD becomes a Galerkin approximation of the Koopman operator since $\langle\mathcal{K}\psi_{k},\psi_{j}\rangle=\langle\psi_{k},\mathcal{K}\psi_{j}\rangle$. Additionally, the formula given by (3.3) is more computationally efficient than the standard symmetric Procrustes approach employed in [13] as it avoids the computation of the SVD of $\mathbf{X}$, as well as a matrix square-root, which can be numerically unstable when $\mathbf{G}$ is ill-conditioned.
Given this convergence result, it is natural to study the convergence of this approximation as the size of the dictionary increases. It is well-known that spectra of Galerkin approximations of operators can suffer from discretization issues such as spectral pollution (spurious eigenvalues), spectral invisibility (missing parts of the spectrum), instabilities, and so forth [45, 46, 47]. These issues occur and, in fact, can worsen as the dictionary’s size increases. Hence, the convergence of the Hermitian DMD algorithm might be difficult to establish. Nevertheless, we will show in Section 4 that the (scalar-valued) spectral measures converge weakly, thus providing theoretical guarantees for Hermitian DMD. One crucial property exploited in the proof is the self-adjointness of the finite-dimensional approximations, as imposed by Hermitian DMD.